# Integrated Information Decomposition Unveils Major Structural Traits of In Silico and In Vitro Neuronal Networks

### v1

@@ V.1 Determination of the relevant information atoms and threshold using in silico neuronal data
In relation to the degree distribution of the reconstruction —see Fig. 4(c)—, the increase in the threshold has the expected effect of reducing the number of links per node. The ground-truth degree distribution is not recovered regardless of the threshold used, indicating that the effective topology never exactly matches the structural one. However, we observe that a higher threshold provides a better definition of a unimodal distribution, as can be seen for the thresholds $z_{th}=2$ and $4$, with the average degree matching the ground-truth for $z_{th}=4$.
@@
Finally, Fig. 4(d) shows the distribution of the effective connection distance. An increase in the threshold decreases the distance of effective connections, showing in fact that the threshold acts as a distance filter, favoring closer elements. Qualitatively, the distributions of the effective connection distance of the effective network agrees well with the ground-truth. However, the observed average value in the effective network is higher than the ground-truth for all the thresholds explored. This result is expected, since the effective interaction between elements usually has a longer range than the structural connection due to the propagation of activity through intermediate nodes. Nevertheless, an additional increase of the threshold favors a closer match between distributions. However, this does not mean that the structural topology exactly agrees with the effective one in any case. Indeed, the apparent match is just an effect of filtering the longest-distance interactions in the effective topology.
@@
Figure 4: Simulated network reconstruction for different thresholds. (First column) Ground truth, structural topology used in simulations. (Others) Functional topology estimated using the unique atom $I_{\partial}^{U_{1}\rightarrow U_{2}}$ and thresholds $z_{th}=1,2,4$ respectively. (a) Network maps, where each dot is a neuron color coded according to the community to which it belongs. (b) Connectivity matrices. Color boxes along the diagonal of the matrices highlight the inferred communities. (c) Degree distribution $P(k_{i})$ of ground-truth (green) and estimations (blue). (d) Connection distance in the ground-truth (green) and effective connection distance in functional reconstructions (blue).
@@ V.2 Application of $\Phi$-ID to in vitro neuronal data
The analysis of the numerical simulations showed that the unique information transfer $I_{\partial}^{U^{1}\rightarrow U^{2}}$ was the atom that contributed the most in reconstructing the ground-truth topology. Thus, we considered this atom to extract key information of the connectivity in a living neuronal network. We considered a threshold of $z_{th}=4$ since, according to the numerical results, the statistics of the effective connectivity (average connectivity and connectivity distances) better matched the ground-truth ones. Experimental data consisted of spontaneous activity recordings of neurons grown on a topographically-modulated substrate akin to the above analyzed simulations. Three experimental repetitions were considered, which were identically prepared but resulted in slightly different networks as they matured.
@@
Figure 5 shows the results of the analysis, as a first observation, we note that activity was not uniform across the culture, and substantially different across repetitions —see Figure 5(a)—, which was caused by fluctuations in the distribution of neurons on the substrate and the intrinsic developmental variability in such a biological, living system. This lack of uniform activity had an effect on network reconstruction. Interestingly, ‘sample 1’ had the most uniform activity across the culture and resulted in an effective connectivity that matched well the imprinted anisotropy of the topographical substrate, i.e., neuronal connectivity reflected the parallel topographical bands that shaped characteristic communities —see Fig. 5(b), left—. In the other samples, although communities clearly existed, they lacked such organization in bands —see Fig. 5(b), center and right—. Nonetheless, as shown in Fig. 5(c), all samples had a similar number of communities and exhibited comparable bulk network properties, with a global efficiency $G_{E}\simeq 0.29$ and a modularity $Q\simeq 0.68$ on average. These results illustrate the difficulty in unveiling key structural traits of living neuronal networks from just activity data, and that effective connectivity cannot be arbitrarily used as a proxy of the structural blueprint of the network.
@@
The results presented here agree with the ones presented in Montala-Flaquer et al. Montalà-Flaquer et al. (2022) where Generalized transfer entropyStetter et al. (2012) was applied. However, here, for the same sample (sample 1), using only the unique transfer atom, we found a larger number of stripe communities (6 here vs 4 in Montalà-Flaquer et al. (2022)) which is closer to the actual pattern, indicating that our reconstruction captures more details of the substrate topography. In addition, we found clear details of the anisotropy for all samples, while in Montala-Flaquer et al. Montalà-Flaquer et al. (2022) this property was only clearly detected in the sample with homogeneous activity (sample 1).
@@
Figure 5(d) compares the in-degree distributions of the experimental repetitions and shows that the three samples share similar distribution shape and average connectivity of $\langle k_{\text{in}}\rangle=10.8\pm 3.1$. This indicates that, despite the clear differences in the spatial organization of communities, the three networks share topological similarities. This is also observed in the distribution of angles between connected neurons —see Figure 5(e)—, where peaks at ${}^{\circ}$ and ${}^{\circ}$ reveal that the anisotropy promoted by the bands imprints a clear directionality in connectivity, i.e., the bands funnel connections along their direction.
@@
To complete the analysis, we also computed the distribution of Euclidean distances between effective connections —see Figure 5(f)—. We observed that samples 1 and 2 had a similar short-range connectivity, with $\langle d_{\text{ec}}\rangle\simeq 1.2$ mm, i.e. neurons effectively connected in a relatively small neighborhood that corresponded to about 20% of the culture’s diameter. Sample 3 had a markedly longer-range connectivity, with $\langle d_{\text{ec}}\rangle\simeq 2.0$ mm (a third of the culture’s diameter). Such a long distance connectivity is clear in the network maps of Figure 5(b), where communities green and maroon horizontally extend much more than the other samples. We note, however, that these communities correspond to areas with relatively poor activity —blue regions of Fig. 5(a)— indicating that they may have evolved somehow isolated from the rest of the culture. This again illustrates how small fluctuations in neuronal density or development of seemingly identical cultures have strong effects on the topological characteristics of mature networks, an aspect intrinsic to biological systems, but very difficult to model theoretically and numerically.
@@
Figure 5: Network reconstruction for three neuronal cultures $6$ mm in diameter with topographical bands. (a) Spatial distribution of spontaneous activity in the cultures, depicting the total number of activations for each ROI in the three different cultures. (b-c) Effective connectivity maps and matrices, derived from the unique information transfer $I_{\partial}^{U^{1}\rightarrow U^{2}}$ with $z_{th}=4$.. Colors identify the inferred communities. (d) Degree distributions $p(k)$ of the effective networks. (e) Distribution of the connection angle $\theta$ in the Euclidean space. (f) Distributions of Euclidean connection distances $d_{ec}$ between ROIs.
@@ VI Discussion
Numerical simulations of living neuronal networks provide a highly valuable environment to explore the capacity of information-theoretic tools to extract major topological features —or ideally the complete ground-truth layout— from spontaneous activity data of the network of interest. Different studies have explored the capacity of models and frameworks to achieve such reconstruction Magrans de Abril, Yoshimoto, and Doya (2018); Banerjee, Chandra, and Ott (2023), and reached the conclusion that a precise reconstruction is in general unattainable given the intrinsic complex nature of a neuronal network, in which cell dynamics, topology and noise may give rise to an almost infinite repertoire of activity patterns. For instance, reconstruction is more difficult when excitatory and inhibitory neurons are considered, or when the spontaneous dynamics of the neuronal network is locked in highly correlated regimes such as network bursts Orlandi et al. (2014). Such a bursting behavior is an innate characteristic of biologically-realistic neuronal networks, both synthetic and living,. Thus, it cannot be easily suppressed and has to be treated as a feature of actual data.
@@
With these considerations in mind, in the present study we showed that $\Phi$-ID can be used to infer aspects of the connectivity of biological neuronal networks. First, we assessed the performance of $\Phi$-ID by comparing the structure inferred from different atoms, in isolation or in different combinations, to the weight matrix derived from a dynamic simulation of a biologically-realistic neuronal network. In this process we were able to identify the relevant information atoms for structure inference. Indeed, transfer entropy (TE) contains, besides the actually information transfer that we are interested in, other information atoms which, for the aim of inferring networks structure from data, contribute as noise. $\Phi$-ID allowed us to decompose the TE measures into different atoms to next consider those atoms that were of relevance. By considering the data from numerical simulations (with known ground-truth topology), we showed that the combination of $I_{\partial}^{U^{1}\to U^{2}}$ and $I_{\partial}^{S\to U^{2}}$, i.e. the unique and synergistic information transfers, respectively, best captured network structure from spontaneous activity.
@@
Simulations also provided a criterion to determine the minimum threshold below which false positive ratio increases abruptly. After obtaining the insights from the simulated data, we defined reasonable threshold values for accepting connections as significant, altogether allowing us to apply the $\Phi$-ID framework to analyze actual biological data obtained from spontaneous activity recordings in neuronal cultures. The results showed that, by using $\Phi$-ID, qualitative aspects of the networks could be extracted, better than by using TE directly. Indeed, $\Phi$-ID was able to reveal key topological features in neuronal cultures associated with imprinted anisotropies, in the form of parallel topographical bands, that promote the development of connections along the direction of the bands.
@@
With the aim to advance in better and more reliable reconstructions of network connectivity from experimental data, it would be interesting to explore in silico the impact of excitatory-inhibitory balance and assess how different dynamic regimes derived from the same ground-truth affect reconstruction. This and other challenges provide an exciting opportunity for the application of new analysis tools such as $\Phi$-ID which could help unveiling some of the secrets of complex systems structure and dynamics.
Finally, the reader should note that the $\Phi$-ID framework is a multivariate extension of Partial Information Decomposition (PID)Barrett (2015) and can also incorporate the methods of pointwise or local information dynamics (LID)Lizier (2013) as is done to calculate the integrated information in cellular automata modelsMediano et al. (2022). In fact, in our application, we are already taking advantage of local information dynamics when, instead of analyzing the whole system with its $N$ elements (neurons or ROIs), we study only the ‘local’ information dynamics of the system by measuring the mutual information atoms of each possible pair $i,j$ of elements. In this sense, the application of $\Phi$-ID to reconstruct the neuronal topology from activity data integrates ideas from both LID and PID, showing the flexibility and generality of the framework.
@@ Acknowledgements.
J.J.T. acknowledges financial support from the Consejería de Transformación Económica, Industria, Conocimiento y Universidades, Junta de Andalucía, Spain and European Regional Development Funds, Ref. P20_00173. This work is also part of the Project of I+D+i, Ref. PID2020-113681GBI00, financed by MICIN/AEI/10.13039/501100011033 and FEDER, Spain, “A way to make Europe”. J.S. and A.M.H. acknowledge financial support from the European Union’s Horizon 2020 research and innovation program under grant agreement No 964977, project NEU-ChiP; by the Ministerio de Ciencia e Innovación, Spain, Grants PID2019-108842GB-C21 and PID2022-137713NB-C22; and by the Generalitat de Catalunya, Grant 2021-SGR-00450. G.M. would like to thank the Programa Nacional de Becas de Postgrados en el Exterior “Don Carlos Antonio López" (BECAL)-Paraguay for the financial support to his doctoral studies in the Physics and Mathematics Program of the University of Granada. G.M also thanks Yasmin May V. for English writing suggestions.

### v2

@@ V.1 Determination of the relevant information atoms and threshold using in silico neuronal data
We also investigated the connectivity in the network through the degree $k$, a general property of graphs that accounts for the number of connections of a node Medaglia, Lynall, and Bassett (2015); Bassett, Zurn, and Gold (2018). Since the reconstructed effective networks are directed, one has to consider in general the incoming connections to a node (in-degree $k_{in}$) and the outgoing ones (out-degree $k_{out}$), with the total connectivity given by $k=k_{in}+k_{out}$. For clarity, Fig. 4(c) shows only the distribution of in-degrees, i.e. the probability to observe a neuron in the network with $k_{in}$ connections. The plots show that an increase in the threshold has the expected effect of reducing the number of links per node. The ground-truth degree distribution is not recovered regardless of the threshold used, indicating that the effective topology never exactly matches the structural one. However, we observe that a higher threshold provides a better definition of a unimodal distribution, as can be seen for the thresholds $z_{\text{th}}=2$ and $4$, with the average degree matching the ground-truth for $z_{\text{th}}=4$.
@@
Finally, Fig. 4(d) shows the distribution of the effective connection distance, i.e. the Euclidean distance between effectively connected neurons. In general, an increase in the threshold decreases the distance, showing in fact that the threshold acts as a distance filter, favoring closer elements. Qualitatively, the distributions of the effective connection distances of the effective network agrees well with the ground-truth. However, the observed average value in the effective network is higher than the ground-truth for all the thresholds explored. This result is expected, since the effective interaction between elements usually has a longer range than the structural connection due to the propagation of activity through intermediate nodes. Nevertheless, an additional increase of the threshold favors a closer match between distributions. However, this does not mean that the structural topology exactly agrees with the effective one in any case. Indeed, the apparent match is just an effect of filtering the longest-distance interactions in the effective topology.
@@
Figure 4: Simulated network reconstruction for different thresholds. (First column) Ground truth, structural topology used in simulations. (Following columns) Effective topology estimated using the unique atom $I_{\partial}^{U_{1}\rightarrow U_{2}}$ and thresholds $z_{\text{th}}=1,2,4$ respectively. (a) Network maps, where each dot is a neuron color coded according to the community to which it belongs. (b) Connectivity matrices. Color boxes along the diagonal of the matrices highlight the inferred communities. Neuron indices have been reordered to increase visibility. (c) Degree distribution of incoming connections $P(k_{in})$ of ground-truth (green) and reconstructed network (blue). (d) Connection distance in the ground-truth (green) and connection distance in reconstructions (blue).
@@ V.2 Application of $\Phi$-ID to in vitro neuronal data
The analysis of the numerical simulations showed that the unique information transfer $I_{\partial}^{U^{1}\rightarrow U^{2}}$ was the atom that contributed the most in reconstructing the ground-truth topology. Thus, we considered this atom to extract key information of the connectivity in a living neuronal network. We considered a threshold of $z_{\text{th}}=4$ since, according to the numerical results, the statistics of the effective connectivity (average connectivity and connectivity distances) better matched the ground-truth ones. We remark that, in general, it is not possible to extract the precise topological organization of a living neuronal network from just activity data, since the neuronal connectivity layout between neurons in combination with intrinsic nonlinear neuronal dynamics and noise can effectually lead to an infinite repertoire of dynamical states. In other words, the ‘inverse problem’ of extracting accurate topological information from activity is not solvable. We therefore aimed here at characterizing the effective connectivity of experimental data and relate it to key features of the studied living networks such as modularity or imprinted anisotropies.
@@
Experimental data consisted of spontaneous activity recordings of neurons grown on a topographically-modulated substrate akin to the above analyzed simulations. Three experimental repetitions were considered, which were identically prepared but resulted in slightly different networks as they matured.
@@
Figure 5 shows the results of the analysis. As a first observation, we note that activity was not uniform across the culture, and substantially different across repetitions —see Figure 5(a)—, which was caused by fluctuations in the distribution of neurons on the substrate and the intrinsic developmental variability in such a biological, living system. This lack of uniform activity had an effect on network reconstruction. Interestingly, ‘sample 1’ had the most uniform activity across the culture and resulted in an effective connectivity that matched well the imprinted anisotropy of the topographical substrate, i.e., neuronal connectivity reflected the parallel topographical bands that shaped characteristic communities —see Fig. 5(b), left—. In the other samples, although communities clearly existed, they lacked such organization in bands —see Fig. 5(b), center and right—. Nonetheless, as shown in Fig. 5(c), all samples had a similar number of communities and exhibited comparable bulk network properties, with a global efficiency $G_{E}\simeq 0.29$ and a modularity $Q\simeq 0.68$ on average. These results illustrate the difficulty in unveiling key structural traits of living neuronal networks from just activity data, and that effective connectivity cannot be arbitrarily used as a proxy of the structural blueprint of the network.
@@
The results presented here agree with the ones presented in Montalà-Flaquer et al. Montalà-Flaquer et al. (2022) where Generalized Transfer EntropyStetter et al. (2012) was applied. However, here, for the same sample (sample 1), using only the unique transfer atom, we found a larger number of stripe communities (6 here vs 4 in Montalà-Flaquer et al. (2022)) which is closer to the actual pattern, indicating that our reconstruction captures more details of the substrate topography. In addition, we found clear details of the anisotropy for all samples, while in Montalà-Flaquer et al. Montalà-Flaquer et al. (2022) this property was only clearly detected in the sample with homogeneous activity (sample 1).
@@
Figure 5(d) compares the distributions of the number of incoming connections $P(k_{\text{in}})$ of the experimental repetitions and shows that the three samples share similar distribution shape and average $\langle k_{\text{in}}\rangle=10.8\pm 3.1$. This indicates that, despite the clear differences in the spatial organization of communities, the three networks share topological similarities. This is also observed in the distribution of angles between connected neurons —see Fig. 5(e)—, where peaks at -90 and 90 reveal that the anisotropy promoted by the bands imprints a clear directionality in connectivity, i.e., the bands funnel connections along their direction.
@@
To complete the analysis, we also computed the distribution of Euclidean distances between effective connections —see Fig. 5(f)—. We observed that samples 1 and 2 had a similar short-range connectivity, with $\langle d_{\text{ec}}\rangle\simeq 1.2$ mm, i.e. neurons effectively connected in a relatively small neighborhood that corresponded to about 20% of the culture’s diameter. Sample 3 had a markedly longer-range connectivity, with $\langle d_{\text{ec}}\rangle\simeq 2.0$ mm (a third of the culture’s diameter). Such a long distance connectivity is clear in the network maps of Fig. 5(b), where communities green and maroon horizontally extend much more than the other samples. We note, however, that these communities correspond to areas with relatively poor activity —blue regions of Fig. 5(a)— indicating that they may have evolved somehow isolated from the rest of the culture. This again illustrates how small fluctuations in neuronal density or development of seemingly identical cultures have strong effects on the topological characteristics of mature networks, an aspect intrinsic to biological systems, but very difficult to model theoretically and numerically.
@@
Figure 5: Network reconstruction for three neuronal cultures $6$ mm in diameter with topographical bands. (a) Spatial distribution of spontaneous activity in the cultures, depicting the total number of activations for each ROI in the three different cultures. (b-c) Effective connectivity maps and matrices, derived from the unique information transfer $I_{\partial}^{U^{1}\rightarrow U^{2}}$ with $z_{\text{th}}=4$. Colors identify the inferred communities. ROI indexes have been reordered to highlight community structure along the diagonal of the matrix. (d) Distribution of incoming connections $P(k_{\text{in}})$ in the reconstructed networks
. (e) Distribution of the connection angle $\theta$ in the Euclidean space. (f) Distributions of effective connection Euclidean distances $d_{ec}$ between ROIs.
@@ VI Discussion
Numerical simulations of living neuronal networks provide a highly valuable environment to explore the capacity of information-theoretic tools to extract major topological features —or ideally the complete ground-truth layout— from spontaneous activity data of the network of interest. Different studies have explored the capacity of models and frameworks to achieve such reconstruction Magrans de Abril, Yoshimoto, and Doya (2018); Banerjee, Chandra, and Ott (2023), and reached the conclusion that a precise reconstruction is in general unattainable given the intrinsic complex nature of a neuronal network, in which cell dynamics, topology and noise may give rise to an almost infinite repertoire of activity patterns. For instance, reconstruction is more difficult when excitatory and inhibitory neurons are considered, or when the spontaneous dynamics of the neuronal network is locked in highly correlated regimes such as network bursts Orlandi et al. (2014). Such a bursting behavior is an innate characteristic of biologically-realistic neuronal networks, both synthetic and living. Thus, it cannot be easily suppressed and has to be treated as a feature of actual data.
An important aspect of Partial Information Decomposition is the need for selecting a redundancy function from many existing candidates Kolchinsky (2022). Our choice of using MMI in the present work was mainly driven by simplicity and low computational cost. However, this redundancy function led to an overestimation of redundancy by considering that the information of the partition with lower mutual information was completely redundant. Thus, MMI sets a lower bound to the synergistic and unique information. For our specific application, we concluded that, among the four $\Phi$-ID atoms that compose the Transfer Entropy measure, the two which are related to redundancy do not substantially contribute to the reconstruction of the network. Therefore, our reconstruction was carried out using the more ‘conservative’ estimation of the relevant atoms. Conservative here means that we are minimizing the estimated information of the relevant atoms. Through this reasoning, we can speculate that redundancy functions which avoid overestimation of redundancy should allow us to obtain even better results. A full exploration of the implications of different redundancy functions to this specific application will be explored in a future work.
@@
With these considerations in mind, in the present study we showed that $\Phi$-ID can be used to infer aspects of the connectivity of biological neuronal networks. First, we assessed the performance of $\Phi$-ID by comparing the structure inferred from different atoms, in isolation or in different combinations, to the weight matrix derived from a dynamic simulation of a biologically-realistic neuronal network. In this process we were able to identify the relevant information atoms for structure inference. Indeed, Transfer Entropy (TE) contains, besides the actually information transfer that we are interested in, other information atoms which, for the aim of inferring networks structure from data, contribute as residuals. $\Phi$-ID allowed us to decompose the TE measures into different atoms to next consider those atoms that were of relevance. By considering the data from numerical simulations (with known ground-truth topology), we showed that the combination of $I_{\partial}^{U^{1}\to U^{2}}$ and $I_{\partial}^{S\to U^{2}}$, i.e. the unique and synergistic information transfers, respectively, best captured network structure from spontaneous activity.
@@
Simulations also provided a criterion to determine the critical threshold below which the false positive rate increases abruptly, and we identified a method to determine this threshold from recorded activity only. After obtaining the insights from the simulated data, we defined reasonable threshold values for accepting connections as significant, altogether allowing us to apply the $\Phi$-ID framework to analyze actual biological data obtained from spontaneous activity recordings in neuronal cultures. The results showed that, by using $\Phi$-ID, qualitative aspects of the networks could be extracted, better than by using TE directly. Indeed, $\Phi$-ID was able to reveal key topological features in neuronal cultures associated with imprinted anisotropies, in the form of parallel topographical bands, that promote the development of connections along the direction of the bands.
@@
With the aim to advance in better and more reliable reconstructions of network connectivity from experimental data, it would be interesting to explore in silico the impact of excitatory-inhibitory balance and assess how different dynamic regimes derived from the same ground-truth affect reconstruction. This and other challenges provide an exciting opportunity for the application of new analysis tools such as $\Phi$-ID which could help unveiling some of the secrets of complex systems’ structure and dynamics.
Finally, the reader should note that the $\Phi$-ID framework is a multivariate extension of Partial Information Decomposition (PID) Barrett (2015) and can also incorporate the methods of pointwise or local information dynamics (LID) Lizier (2013) as is done to calculate the integrated information in cellular automata models Mediano et al. (2022). In fact, in our application, we are already taking advantage of local information dynamics when, instead of analyzing the whole system with its $N$ elements (neurons or ROIs), we study only the ‘local’ information dynamics of the system by measuring the mutual information atoms of each possible pair $i,j$ of elements. In this sense, the application of $\Phi$-ID to reconstruct the neuronal topology from activity data integrates ideas from both LID and PID, showing the flexibility and generality of the framework. In our particular case, we applied the framework to boolean time series, which simplify the estimation of probability distributions. When working with continuous time series, different strategies can be applied, e.g. the use of entropy estimators such as the ones based on k-neighbors methods  Kraskov, Stögbauer, and Grassberger (2004) or linear-gaussian estimators as the ones applied for blood-oxygen-level-dependent (BOLD) signals in Luppi et al. (2022) to compute mutual information. However, we should bear in mind that there are no estimators free of biases and limitations Xiong, Faes, and Ivanov (2017), specially when dealing with limited continuous variables such as the ones obtained from real complex systems. Nonetheless, these limitations are common to any information theory measure, not only to the framework applied in this work.
@@ Acknowledgements.
J.J.T. acknowledges financial support from the Consejería de Transformación Económica, Industria, Conocimiento y Universidades, Junta de Andalucía, Spain and European Regional Development Funds, Ref. P20_00173. This work is also part of the Project of I+D+i, Ref. PID2020-113681GBI00, financed by MICIN/AEI/10.13039/501100011033 and FEDER, Spain, “A way to make Europe”. J.S. and A.M.H. acknowledge financial support from the European Union’s Horizon 2020 research and innovation program under grant agreement No 964877, project NEU-ChiP; by the Ministerio de Ciencia e Innovación, Spain, Grants PID2019-108842GB-C21 and PID2022-137713NB-C22; and by the Generalitat de Catalunya, Grant 2021-SGR-00450. G.M. would like to thank the Programa Nacional de Becas de Postgrados en el Exterior “Don Carlos Antonio López" (BECAL)-Paraguay for the financial support to his doctoral studies in the Physics and Mathematics Program of the University of Granada. G.M also thanks Yasmin May V. for English writing suggestions.